# Enhancing Vision-Language Models with Scene Graphs for Traffic Accident Understanding

### v2

@@ V-B1 Pre-trained SGE with Alignment
Regarding the case when the model is pre-trained on the DoTA dataset, the model actually has a slightly better performance without undergoing the alignment training (Column 5) versus with alignment (Column 4). This seems to suggest that the model is not learning further useful information during alignment training. One possible explanation for this is that the data domains of the frozen encoders are not similar enough to DoTA for any positive transfer of data domain knowledge to occur Zhang et al. (2023). Another possibility has to do with the alignment training procedure, which is based on CLIP’s training algorithm and therefore relies heavily on using a large batch size at each step Radford et al. (2021). This is because the algorithm performs contrastive training on each of the three pairs of embedding spaces (vision-graph, text-graph, and vision-text) for each batch of traffic accident examples. However, given the limited size of the DoTA dataset and compute constraints, our experiments in Table II use a batch size of 32 and are limited to 10 epochs of training. We hypothesize that a larger batch size and/or number of training epochs would increase the accuracy gain of the training alignment step. To test this hypothesis, we carried out some additional experiments, the results of which are shown in Table III, and indicate that both increasing batch size and the number of epochs give similar results to the “SGE without Alignment” case shown in Table II (Column 5) rather than worse results (Columns 2-4 of Table II). It is possible that further training of the larger batch sizes would have improved these results beyond the values in Column 5.
TABLE III: Alignment training with different hyperparameter configurations, using a SGE pre-trained on DoTA. “e” is number of epochs, “bs” is batch size for the alignment training. Column 4 in Table II is the same as Column 1 in this table.
@@ V-B2 No Pre-trained SGE with Alignment
TABLE IV: Training for alignment from scratch with different configurations. Column 2 in Table II is the same as Column 1 in this table.
@@ VI Conclusion and Future Work
There are several areas where future work can be done in this task. Firstly, although three modalities are used, little exploration has been conducted on the language modality. In principle, captions for each video should be derived from the video itself, as the SGs are. This would allow for the captions to be fed into the model to fine-tune the classifier head and run inference with a greater signal coming from the language modality. Next, although they provide some benefit for the model, the generated SGs have a relatively limited structure: only categorical distance relations between a vehicle and objects in its surroundings are generated, along with mappings to a fixed set of three traffic lanes (left, middle, and right). Enabling a semantic extension of the generated SGs may enhance the signal obtained from this modality. Similarly, by modeling relations between all objects in the scene, this pipeline could be expanded for the non-ego case as well, and perhaps could be used on more accident classes in the DoTA dataset or similar use cases. Regarding the dataset, although we attempt to resolve the data imbalance by focusing on the balanced accuracy metric, it would be interesting to explore data augmentation techniques as well. Additionally, the MRGCN-based architecture of the SGE was taken directly from Yu et al. (2020), but perhaps it can be further improved by modifying either its spatial or temporal modeling components. Given the modular design of our system, another fascinating avenue of exploration would be to incorporate an additional modality (e.g., audio) to include in the alignment step before fusing the embeddings and fine tuning the classification head. Experiments with different modality fusion methods as well as classification heads may also lead to improvements.
@@ Appendix A Generated Scene Graph Examples
Figure 2: A scene graph generated using the rs2v Malawade et al. [2021] tool. Starting from a video frame (Raw Image), the SGG first detects objects in the scene (Object Detection Image) and generates the BEV (Bird’s Eye Image) before creating the scene graph representation (SceneGraph Image). This scene graph shows the ego car relative to two other vehicles, one categorized as in the left lane, the other right. The closer vehicle is recognized as being near collision (with the edge attribute “near_coll”), whereas the farther vehicle is registered in the scene graph as simply being “visible”.

### v3

@@ V-B1 Pre-trained SGE with Alignment
Regarding the case when the model is pre-trained on the DoTA dataset, the model actually has a slightly better performance without undergoing the alignment training (Column 5) versus with alignment (Column 4). This seems to suggest that the model is not learning further useful information during alignment training. One possible explanation for this is that the data domains of the frozen encoders are not similar enough to DoTA for any positive transfer of data domain knowledge to occur [35]. Another possibility has to do with the alignment training procedure, which is based on CLIP’s training algorithm and therefore relies heavily on using a large batch size at each step [10]. This is because the algorithm performs contrastive training on each of the three pairs of embedding spaces (vision-graph, text-graph, and vision-text) for each batch of traffic accident examples. However, given the limited size of the DoTA dataset and compute constraints, our experiments in Table II use a batch size of 32 and are limited to 10 epochs of training. We hypothesize that a larger batch size and/or number of training epochs would increase the accuracy gain of the training alignment step. To test this hypothesis, we carried out some additional experiments, the results of which are shown in Table III, and indicate that both increasing batch size and the number of epochs give similar results to the “SGE without Alignment” case shown in Table II (Column 5) rather than worse results (Columns 2-4 of Table II). It is possible that further training of the larger batch sizes would have improved these results beyond the values in Column 5.
TABLE III: Alignment training with different hyperparameter configurations, using a SGE pre-trained on DoTA. “e” is number of epochs, “bs” is batch size for the alignment training. Column 4 in Table II is the same as Column 1 in this table
@@ V-B2 No Pre-trained SGE with Alignment
TABLE IV: Training for alignment from scratch with different configurations. Column 2 in Table II is the same as Column 1 in this table
@@ VI Conclusion and Future Work
There are several areas where future work can be done in this task. Firstly, although three modalities are used, little exploration has been conducted on the language modality. In principle, captions for each video should be derived from the video itself, as the SGs are. This would allow for the captions to be fed into the model to fine-tune the classifier head and run inference with a greater signal coming from the language modality. Next, although they provide some benefit for the model, the generated SGs have a relatively limited structure: only categorical distance relations between a vehicle and objects in its surroundings are generated, along with mappings to a fixed set of three traffic lanes (left, middle, and right). Enabling a semantic extension of the generated SGs may enhance the signal obtained from this modality. Similarly, by modeling relations between all objects in the scene, this pipeline could be expanded for the non-ego case as well, and perhaps could be used on more accident classes in the DoTA dataset or similar use cases. Regarding the dataset, although we attempt to resolve the data imbalance by focusing on the balanced accuracy metric, it would be interesting to explore data augmentation techniques as well. Additionally, the MRGCN-based architecture of the SGE was taken directly from [12], but perhaps it can be further improved by modifying either its spatial or temporal modeling components. Given the modular design of our system, another fascinating avenue of exploration would be to incorporate an additional modality (e.g., audio) to include in the alignment step before fusing the embeddings and fine tuning the classification head. Experiments with different modality fusion methods as well as classification heads may also lead to improvements.
@@ Appendix A Generated Scene Graph Examples
Figure 2: A scene graph generated using the rs2v [8] tool. Starting from a video frame (Raw Image), the SGG first detects objects in the scene (Object Detection Image) and generates the BEV (Bird’s Eye Image) before creating the scene graph representation (SceneGraph Image). This scene graph shows the ego car relative to two other vehicles, one categorized as in the left lane, the other right. The closer vehicle is recognized as being near collision (with the edge attribute “near_coll”), whereas the farther vehicle is registered in the scene graph as simply being “visible”.